1/2

Дмитрий Николаевич Садовников,  Дмитрий Николаевич Садовников,
Sten’ka Razin  Stenka Rasin
tradukita de Tr. Jacob Robbins  tradukita de N. N. 21

1. From beyond the wooded island To the river wide and free Proudly sailed the arrow-breasted Ships of Cossack yeomanry.

In den Wellen hinter Inseln Ziehen Kähne malerisch, - |: Fangen leis an aufzuwachen, Bang ist jedes Angesicht. :|

2. On the first is Stenka Razin With a princess at his side, |: Drunken, holds a marriage revel With his beautiful young bride. :| Stenka Rasin vorn als erster, Selig in der Trunkenheit, |: Hält im Arme die Prinzessin, Die er eben erst befreit. :|

3. But behind them rose a whisper, “He has left his sword to woo; |: One short night, and Stenka Razin Has become a woman too!” :|

.................................................. .................................................. .................................................. ..................................................

4. Stenka Razin hears the jeering Of his discontented band, |: And the lovely Persian princess He has circled with his hand. :|

.................................................. .................................................. .................................................. ..................................................

5. His black brows have come together As the waves of anger rise, |: And the blood comes rushing swiftly To his piercing, jet-black eyes. :|

.................................................. .................................................. .................................................. ..................................................

6. “I will give you all you ask for, Life and heart, and head and hand,” |: Echo rolls the pealing thunder Of his voice across the land. :|

.................................................. .................................................. .................................................. ..................................................

7. “Volga, Volga, mother Volga, Deep and wide beneath the sun, |: You have never seen a present From the Cossack of the Don. :| Plötzlich tönt ein dumpf Gemurre: Er verrät uns um ein Weib, |: All der Seinen Glück vergißt er Um geringen Zeitvertreib. :|

8. And that peace might rule as always All my free-born men and brave, |: Volga, Volga, mother Volga, Volga, make this girl a grave.” :| Wolga, Wolga, Mutter Wolga, Ruhig fließest du dahin, |: Ahnst nicht, was ein Donkosake Fürchterliches hat im Sinn. :|

…  …

2/2

9. With a sudden, mighty movement, Razin lifts the beauty high, |: And he casts her where the waters Of the Volga move and sigh. :|

.................................................. .................................................. .................................................. ..................................................

10. Now a silence like the grave sinks To all those who stand to see, |: And the battle-hardened Cossacks Sink to weep on bended knee. :| Vorn als erster Stenka Rasin Hebt das Weib in wilder Wut, - |: Wolga, Wolga, nimm ein Opfer! Und er wirft sie in die Flut. :|

11. “Dance, you fool, and men, make merry! What has got into your eyes? |: Let us thunder out a chanty Of a place where beauty lies.” :| Und er sieht sie untergehen, Hört noch ihren Jammerschrei: |: Stirb als Opfer meiner Treue! Stenka Rasin, er ist frei. :|

12. From beyond the wooded island To the river wide and free, |: Proudly sail the arrow-breasted Ships of Cossack yeomanry. :| Und die Kähne ziehen weiter, Und die Kähne ziehen fort, |: Und die Wolga fließet weiter Über diesen Schreckensort.

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Germanan de N. N. 21.

Arg-478-1038 (2010-03-09 15:31:43)

Mi, Manfred Retzlaff, trovis tiun ĉi germanigon de la kanto en http: // www. musicanet. org/ robokopp/ Lieder/ stenkara. html . Latinskriba aliskribo kaj laŭvorta germanigo troveblas en http: // www. kaikracht. de/ balalaika/ songs/ izza_bal. htm . Vidu ankaŭ la vikipedian retejon http: // de. wikipedia. org/ wiki/ Stenka_Rasin .